# Advancing Depth Anything Model for Unsupervised Monocular Depth Estimation in Endoscopy

### v2

@@ I INTRODUCTION
Figure 2: Network architecture diagram. The encoder in the depth network contains 12 transformer modules, each with an RVLoRA module. Res-DSC modules are added after the 3rd, 6th, 9th, and 12th layers.
@@
There are numerous strategies for fine-tuning foundation models, with Low-Rank Adaptation (LoRA) being one of the most commonly used methods [14]. However, LoRA does not always perform optimally and some variations show better performance in certain tasks. To address this, we explore a novel fine-tuning strategy aimed at enhancing performance without increasing the number of parameters in LoRA. Inspired by recent studies that highlight the surprising effectiveness of models using random weights and projections [15, 16], we propose a low-rank adaptation method based on random vectors (RVLoRA) that enhances model performance without adding additional trainable parameters. Additionally, transformer models tend to focus more on low-frequency information, resulting in insufficient ability to extract texture details and local features [17, 18]. To address this, we propose a residual block based on depthwise separable convolution (Res-DSC) to achieve higher performance by enhancing locality.
@@ III-A Image Intrinsic-Based UMDE
This framework relies on the assumption of photometric consistency, which no longer holds in endoscopy. IID-SfMLearner [6] employs intrinsic image decomposition to tackle the challenge, making it the most advanced framework for monocular endoscopic depth estimation. IID-SfMLearner additionally introduces the intrinsic decomposition module and synthetic reconstruction module to perform image decomposition and refine the decomposition results. This allows the method to leverage the inherent reflectance consistency of images, providing better supervision during network training.
@@ III-B1 Low-rank Adaptation Based on Random Vectors (RVLoRA)
Figure 4: The detailed structure of the Res-DSC module.
@@ III-C Total Training Framework
SfMLearner [4] CVPR’17 0.086$\pm$0.037 1.021$\pm$1.100 7.553$\pm$5.110 0.121$\pm$0.060 0.925$\pm$0.081 31.60 31.60
Monodepth2 [5] ICCV’19 0.066$\pm$0.025 0.577$\pm$0.482 5.781$\pm$3.174 0.093$\pm$0.036 0.961$\pm$0.043 14.84 14.84
Endo-SfMLearner [20] MedIA’21 0.068$\pm$0.029 0.679$\pm$0.709 6.227$\pm$4.137 0.098$\pm$0.047 0.955$\pm$0.054 14.84 14.84
AF-SfMLearner [7] MedIA’22 0.060$\pm$0.024 0.443$\pm$0.347 4.964$\pm$2.348 0.082$\pm$0.032 0.973$\pm$0.037 14.84 14.84
IID-SfMLearner [6] JBHI’24 0.058$\pm$0.026 0.435$\pm$0.375 4.820$\pm$2.583 0.080$\pm$0.035 0.969$\pm$0.043 14.84 14.84
Depth Anything [12] CVPR’24 0.086$\pm$0.039 0.927$\pm$0.875 6.957$\pm$3.817 0.112$\pm$0.049 0.929$\pm$0.089 97.50 97.50
EndoDAC [13] MICCAI’24 0.051$\pm$0.022 0.355$\pm$0.334 4.442$\pm$2.509 0.073$\pm$0.032 0.979$\pm$0.032 99.09 1.66
Ours - 0.048$\pm$0.021 0.315$\pm$0.297 4.172$\pm$2.261 0.068$\pm$0.030 0.982$\pm$0.027 98.80 1.38
@@ IV-C Comparison Experiments
We compared our method with seven competitive unsupervised depth estimation methods: SfmLearner [4], Monodepth2 [5], Endo-SfmLearner [20], AF-SfmLearner [7], IID-SfmLearner [6], Depth Anything [12], and EndoDAC [13]. Among these, EndoDAC is currently the state-of-the-art method on the SCARED dataset. The first three methods are reproduced from scratch following the code provided by the authors, and the last four methods are tested using the optimal model provided by the authors. Table II presents the results of the comparative experiments.
SfMLearner and Monodepth2, which are designed for natural scenes, perform poorly on endoscopic images. While Endo-SfMLearner, AF-SfMLearner, and IID-SfMLearner have introduced algorithmic improvements tailored to the unique characteristics of endoscopic images, their performance gains remain limited. Depth Anything Model, the latest foundation model for depth estimation, also falls short due to the absence of medical images in its training datasets. EndoDAC, which similarly fine-tunes the Depth Anything Model, has achieved significant improvements over earlier methods. In comparison, our method outperforms all others across various metrics. Notably, our approach has the fewest trainable parameters, totaling only 1.38M, which accounts for just 1.4% of the overall parameters.
@@ IV-D Ablation Study
In our method, the Res-DSC modules are evenly distributed across the 12 transformer modules of the encoder. For instance, when using three Res-DSC modules, they are placed at the 4th, 8th, and 12th layers. To determine the optimal number of Res-DSC modules, we conducted a series of ablation experiments. The results, shown in Table VI, indicate that using four Res-DSC modules achieves the best performance while maintaining a relatively low parameter count.
@@
TABLE VI: The ablation study on the number of Res-DSC modules
@@
TABLE VII: The results of pose estimation.
@@ IV-E Pose Estimation
In unsupervised monocular depth estimation, camera pose estimation and depth estimation are jointly learned tasks, where the accuracy of the pose network’s estimations can also indirectly reflect the precision of the depth estimation. Therefore, we also evaluated the performance of the pose network on the SCARED dataset. We assessed three video sequences using the Absolute Trajectory Error (ATE) metric, comparing our results only with the most competitive method, EndoDAC [13]. The quantitative results are presented in Table VII, where our method outperforms the compared method across the three sequences. The visual results, shown in Fig. 7, demonstrate reduced trajectory drift and a closer alignment with the ground truth for our method. This further validates the effectiveness of our depth estimation approach.

### v3

@@ I INTRODUCTION
Figure 2: Network architecture diagram. The DepthNet is used to predict the depth information. The PoseNet is used to estimate the camera’s pose. The intrinsic decomposition module is used to decompose the images into reflectance and shading components. The synthesis reconstruction module is used to reconstruct the target image. The training method is the same with the IID-SfMLearner [6]. The encoder in the depth network contains 12 transformer modules, each with an RVLoRA module. Res-DSC modules are added after the 3rd, 6th, 9th, and 12th layers.
@@
There are numerous strategies for fine-tuning foundation models, with Low-Rank Adaptation (LoRA) being one of the most commonly used methods [14]. However, LoRA does not always perform optimally and some variations show better performance in certain tasks [15]. To address this, we explore a novel fine-tuning strategy aimed at enhancing performance without increasing the number of parameters in LoRA. Inspired by recent studies that highlight the surprising effectiveness of models using random weights and projections [16, 17], we propose a low-rank adaptation method based on random vectors (RVLoRA) that enhances model performance without adding additional trainable parameters. Additionally, transformer models tend to focus more on low-frequency information, resulting in insufficient ability to extract texture details and local features [18, 19]. To address this, we propose a residual block based on depthwise separable convolution (Res-DSC) to achieve higher performance by enhancing locality.
@@ III-A Image Intrinsic-Based UMDE
This framework relies on the assumption of photometric consistency, which no longer holds in endoscopy. IID-SfMLearner [6] employs intrinsic image decomposition to tackle the challenge, making it the most advanced framework for monocular endoscopic depth estimation. IID-SfMLearner additionally introduces the intrinsic decomposition module and synthetic reconstruction module to decompose the image into reflectance and shading images. This allows the method to leverage the inherent reflectance consistency of images, providing better supervision during network training.
@@ III-B1 Low-rank Adaptation Based on Random Vectors (RVLoRA)
Figure 4: The detailed structure of the Res-DSC module. Res-DSC contains a depthwise separable convolution, a pointwise separable convolution, and the residual connection.
@@ III-C Total Training Framework
SfMLearner [4] 0.086$\pm$0.037 1.021$\pm$1.100 7.553$\pm$5.110 0.121$\pm$0.060 0.925$\pm$0.081 31.60 31.60
Monodepth2 [5] 0.066$\pm$0.025 0.577$\pm$0.482 5.781$\pm$3.174 0.093$\pm$0.036 0.961$\pm$0.043 14.84 14.84
Endo-SfMLearner [21] 0.068$\pm$0.029 0.679$\pm$0.709 6.227$\pm$4.137 0.098$\pm$0.047 0.955$\pm$0.054 14.84 14.84
AF-SfMLearner [7] 0.060$\pm$0.024 0.443$\pm$0.347 4.964$\pm$2.348 0.082$\pm$0.032 0.973$\pm$0.037 14.84 14.84
IID-SfMLearner [6] 0.058$\pm$0.026 0.435$\pm$0.375 4.820$\pm$2.583 0.080$\pm$0.035 0.969$\pm$0.043 14.84 14.84
Depth Anything [12] 0.086$\pm$0.039 0.927$\pm$0.875 6.957$\pm$3.817 0.112$\pm$0.049 0.929$\pm$0.089 97.50 97.50
EndoDAC [13] 0.051$\pm$0.022 0.355$\pm$0.334 4.442$\pm$2.509 0.073$\pm$0.032 0.979$\pm$0.032 99.09 1.66
Ours 0.048$\pm$0.021 0.315$\pm$0.297 4.172$\pm$2.261 0.068$\pm$0.030 0.982$\pm$0.027 98.80 1.38
@@ IV-C Comparison Experiments
We compared our method with seven competitive unsupervised depth estimation methods: SfmLearner [4], Monodepth2 [5], Endo-SfmLearner [21], AF-SfmLearner [7], IID-SfmLearner [6], Depth Anything [12], and EndoDAC [13]. Among these, EndoDAC is currently the state-of-the-art method on the SCARED dataset. The first three methods are reproduced from scratch following the code provided by the authors, and the last four methods are tested using the optimal model provided by the authors. Table II presents the experimental results and parameter numbers of the comparative experiments.
SfMLearner and Monodepth2, which are designed for natural scenes, perform poorly on endoscopic images. While Endo-SfMLearner, AF-SfMLearner, and IID-SfMLearner have introduced algorithmic improvements tailored to the unique characteristics of endoscopic images, their performance gains remain limited. Depth Anything Model, the latest foundation model for depth estimation, also falls short due to the absence of medical images in its training datasets. EndoDAC, which similarly fine-tunes the Depth Anything Model, has achieved significant improvements over earlier methods. Compared to EndoDAC, our method improves the fine-tuning approach and reduces the number of trainable parameters. In comparison, our method outperforms all others across various metrics. Notably, our approach has the fewest trainable parameters, totaling only 1.38M, which accounts for just 1.4% of the overall parameters.
@@ IV-D Ablation Study
Compared to other initialization methods of the frozen parameters, Table VI shows the effectiveness of Kaiming uniform initialization.
In our method, the Res-DSC modules are evenly distributed across the 12 transformer modules of the encoder. For instance, when using three Res-DSC modules, they are placed at the 4th, 8th, and 12th layers. To determine the optimal number of Res-DSC modules, we conducted a series of ablation experiments. The results, shown in Table VII, indicate that using four Res-DSC modules achieves the best performance while maintaining a relatively low parameter count.
@@
TABLE VI: The ablation study on the vector initialization.
[…]
Uniform 0.050 0.333 4.328 0.070 0.983
Kaiming Normal 0.050 0.336 4.201 0.071 0.976
Kaiming Uniform 0.048 0.315 4.172 0.068 0.982
TABLE VII: The ablation study on the number of Res-DSC modules.
@@
TABLE VIII: The results of pose estimation.
@@ IV-E Pose Estimation
In unsupervised monocular depth estimation, camera pose estimation and depth estimation are jointly learned tasks, where the accuracy of the pose network’s estimations can also indirectly reflect the precision of the depth estimation. Therefore, we also evaluated the performance of the pose network on the SCARED dataset. We assessed three video sequences using the Absolute Trajectory Error (ATE) metric, comparing our results only with the most competitive method, EndoDAC [13]. The quantitative results are presented in Table VIII, where our method outperforms the compared method across the three sequences. The visual results, shown in Fig. 7, demonstrate reduced trajectory drift and a closer alignment with the ground truth for our method. This further validates the effectiveness of our depth estimation approach.